### v2

# Large Language Models as Quasi-crystals: Coherence Without Repetition in Generative Text

Abstract
This essay proposes an interpretive analogy between large language models (LLMs) and quasicrystals, systems that exhibit global coherence without periodic repetition, generated through local constraints. While LLMs are typically evaluated in terms of predictive accuracy, factuality, or alignment, this structural perspective suggests that one of their most characteristic behaviors is the production of internally resonant linguistic patterns. Drawing on the history of quasicrystals, which forced a redefinition of structural order in physical systems, the analogy highlights an alternative mode of coherence in generative language: constraint-based organization without repetition or symbolic intent. Rather than viewing LLMs as imperfect agents or stochastic approximators, we suggest understanding them as generators of quasi-structured outputs. This framing complements existing evaluation paradigms by foregrounding formal coherence and pattern as interpretable features of model behavior. While the analogy has limits, it offers a conceptual tool for exploring how coherence might arise and be assessed in systems where meaning is emergent, partial, or inaccessible. In support of this perspective, we draw on philosophy of science and language, including model-based accounts of scientific representation, structural realism, and inferentialist views of meaning. We further propose the notion of structural evaluation: a mode of assessment that examines how well outputs propagate constraint, variation, and order across spans of generated text. This essay aims to reframe the current discussion around large language models, not by rejecting existing methods, but by suggesting an additional axis of interpretation grounded in structure rather than semantics.
I Introduction
Large language models (LLMs) are statistical systems trained on massive corpora of human text to generate coherent language outputs. Developed through deep learning architectures such as the Transformer, [1] these models operate by predicting the most probable next token in a sequence based on prior context. A token is a basic unit of text (typically a word, subword, or punctuation mark) used to segment language for computational modeling. Each new token is selected in relation to all those that came before, allowing the model to generate fluent sequences without predefined grammar or semantic rules. Since the release of GPT-2 in 2019 [2] and especially GPT-3 in 2020, [3] which helped inaugurate the class of foundation models, [4] LLMs have shown an ability to produce language that is locally fluent and contextually plausible across a wide range of domains. These outputs are not the result of symbolic reasoning or understanding, but of pattern recognition applied at large scale. The resulting text often exhibits internal consistency, thematic alignment, and the appearance of intention, even when none is explicitly encoded. [5]
Quasicrystals are a class of solid materials that exhibit long-range order without periodic repetition. First identified in 1982 by Dan Shechtman, [6] their discovery challenged the prevailing assumption that atomic arrangements in crystals must repeat regularly in space. Quasicrystals display symmetry, such as fivefold rotations, that are forbidden in classical crystallography, yet their structure is neither random nor disordered. Instead, they possess deterministic but aperiodic arrangements, as first modeled by Levine and Steinhardt using tiling patterns such as those of Roger Penrose. [7] The recognition of quasicrystals as a legitimate form of structural order expanded the conceptual boundaries of solid-state chemistry and materials science, opening new directions in the study of symmetry, entropy, and self-assembly. [8]
We propose that the structure of LLM-generated text exhibits properties analogous to those of quasicrystals. The analogy applies specifically to the output: sequences of language produced through token-level prediction, not the internal architecture of the models themselves. Like quasicrystalline matter, this output displays local coherence without global repetition. It is neither disordered nor strictly periodic, but unfolds according to statistical constraints that produce a recognizably structured, yet non-repeating, surface. [10] Just as quasicrystals forced a redefinition of what constitutes order in physical systems, LLMs challenge inherited assumptions about coherence in language. Other systems, such as cellular automata, also generate complex behavior from simple local rules and have been widely studied as models of emergent structure. [9] However, quasicrystals are distinctive in that they introduced a formally recognized category of long-range aperiodic order in the physical sciences. This makes them especially apt as analogues for linguistic outputs that exhibit coherence without explicit rules or symbolic intent. In both cases, structured form emerges not from central planning, but from distributed interactions that propagate constraint across a generative space. [14]
This analogy offers more than surface resemblance, it can clarify why LLM outputs exhibit a surprising degree of coherence without explicit planning, and why they resist both traditional linguistic analysis and formal logical modeling. While terms like “sense” and “coherence” remain philosophically open, we use them here in a structural sense: to describe the appearance of internal consistency, recurrence, and thematic alignment across spans of generated text. This perspective offers a middle ground between dismissive accounts that see LLMs as stochastic parrots [5] and overly anthropomorphic ones that frame them as emergent minds. [14] By emphasizing their capacity to generate constraint-based linguistic forms, we gain a language for describing the generative space LLMs inhabit, one that accounts for novelty without invoking consciousness. This shift may support more precise discussions of authorship, creativity, and trust in machine-generated language, [4] as well as inform future model design across domains that demand both structure and novelty.
The analogy we draw is not intended as an exact isomorphism between physical and linguistic systems, but as a model-based framework in the sense described by Hesse [11] and Suárez [12]: a heuristic device that preserves relevant structure and supports new modes of inquiry. In this case, the quasicrystal model provides a way to conceptualize LLM outputs as emergent structures governed by local constraint rather than global intent. As with many productive scientific analogies, the value lies not in literal correspondence but in enabling reframing and making visible aspects of coherence and generation that standard evaluations might overlook.
The remainder of this essay unfolds in three parts. We first summarize the relevant properties of quasicrystals and LLMs, with particular attention to how structure arises in each system from local constraints. We then develop the analogy between them, highlighting how coherence and form emerge without repetition or centralized control. Building on this, we outline practical implications for evaluating and training language models, suggesting that their behavior may be better understood in structural rather than semantic terms. Finally, we conclude with a reflection on how this analogy reframes questions of coherence, meaning, and form in generative systems more broadly.
II Order without periodicity
Quasicrystals display a form of structural order that emerges without translational periodicity. Unlike conventional crystals, which repeat unit cells in regular three-dimensional arrays, quasicrystals arrange atoms according to deterministic rules that never produce a repeated pattern. This leads to symmetries considered forbidden in classical crystallography, such as fivefold rotations. [6] Their diffraction patterns contain sharp Bragg peaks, confirming long-range order, but they cannot be indexed using the lattice parameters of periodic crystals. [7]
To describe these materials, scientists have relied on mathematical tiling schemes, such as Penrose tiling in two dimensions and Ammann tilings in three. [8] These models break space into basic geometric shapes—tiles—that must fit together according to specific local rules. Unlike in periodic crystals, these tilings never repeat exactly, yet they still generate long-range order. The local constraints propagate outward to form globally structured, non-periodic arrangements. The resulting patterns maintain coherence across large distances without regular repetition, revealing a form of organization that sits between randomness and strict regularity. The recognition of this category required a broadening of how structural order is defined, both in physical systems and in the conceptual language of symmetry. [8, 7]
Large language models generate text by predicting the most likely next token in a sequence, given a preceding context. This process is not governed by rules of syntax or meaning in the traditional sense, but by probabilistic correlations learned from large datasets of human language. Each output is the result of locally conditioned prediction: the model selects a token based on the statistical structure of the tokens before it, with no explicit awareness of global structure or intent (though in practice, the attention mechanism grants access to the full preceding context). Yet, the generated language often appears coherent across multiple sentences or even paragraphs, suggesting a kind of emergent organization. [10]
This emergent structure arises from the scale and architecture of modern LLMs. With billions of parameters and attention mechanisms that capture relationships between distant tokens, [1] these models encode a diffuse statistical map of language use. [4] Coherence is not programmed, but surfaced. It is a byproduct of local decisions cascading under constraint. As with tiling schemes in quasicrystals, small, local rules can give rise to patterns that feel organized, readable, and meaningful over extended spans, even in the absence of explicit design. The result is a form of language that is neither scripted nor random, but ordered in ways that resist simple repetition or deterministic unfolding.
Despite operating in different domains, material and linguistic, quasicrystals and large language models both generate structured patterns through systems of local constraint. [7] In quasicrystals, tiling rules enforce spatial relationships that propagate non-repeated, ordered patterns. In LLMs, probabilistic models of token relationships produce text that unfolds in coherent but non-deterministic ways. Neither system relies on a global plan or repeated unit, yet both give rise to forms that are internally consistent and often perceived as intentional or meaningful. This resemblance is more than superficial: it suggests a shared structural logic that may help explain the growing sense that LLM-generated language “makes sense” even when it is not grounded in symbolic understanding.
This analogy does not imply equivalence between material and linguistic order, but instead highlights a zone of structural similarity where pattern emerges without repetition and coherence does not require blueprinting. Considering LLM outputs as linguistic analogues to quasicrystalline structure offers a new framework for interpreting their generative behavior not as noise, nor mimicry, but as a distinct form of constrained emergence. This framing may inform how we evaluate, design, and trust generative systems, not by replacing traditional metrics of logic or narrative, but by supplementing them with attention to structural coherence, propagation, and form.
III Toward Structural Evaluation
Recent claims about LLMs’ capacity for reasoning, planning, or logic [13, 14] might appear to challenge this framing. However, we suggest that such behaviors are better understood as high-level manifestations of statistical coherence, rather than evidence of symbolic reasoning in the traditional sense. As others have argued, LLMs generate plausible approximations of reasoning by operating over patterns learned from text, not by manipulating abstract concepts or representations with intent. [5, 15] Our analogy focuses on this structural dimension: the model’s ability to produce outputs that exhibit internal consistency and pattern without invoking an explicit logic or plan.
This structural perspective also exposes a limitation in how large language models are currently evaluated. These systems are typically assessed through functional lenses: accuracy in downstream tasks, consistency with human preferences, or alignment with external truths. [4, 16] While these metrics are valuable, they often treat coherence as a secondary byproduct of optimization rather than as a central feature. If we take seriously the analogy between LLMs and quasicrystals of language, then coherence is not incidental but fundamental. In this view, an LLM is not merely a predictor of plausible continuations, but a generator of constrained structure. Its success lies not only in what it says, but in how its internal form sustains tension, variation, and order over time.
This reframing also suggests new directions for evaluation. If LLMs are generative systems whose characteristic behavior is the propagation of internal structure, then we might assess their outputs in terms of formal properties rather than semantic correctness. Structural features such as local-to-global coherence, the recurrence of non-identical motifs, or the statistical rhythm of variation become relevant indicators. Methods inspired by signal processing, including entropy mapping, motif extraction, or Fourier-like decomposition, could surface organizational patterns that token-level metrics fail to detect. For instance, an entropy heat-map of a 1000-token completion could reveal zones of high and low unpredictability, indicating shifts in internal structure, theme, or register. Recurrent but non-identical motifs might suggest narrative pacing or topical development. Such features would allow for diagnostics that go beyond correctness to reflect how well a model sustains coherent variation over time. In this light, structural resonance, the degree to which an output holds together through internal patterning, might become a meaningful evaluative target, particularly in domains such as storytelling, dialogue, or abstraction, where formal coherence is often more valuable than semantic precision.
Seen through this lens, structural evaluation becomes a complementary mode of assessment. It does not replace truthfulness, safety, or usefulness as metrics, but foregrounds a different kind of judgment: whether the output of a model is internally consistent, non-redundant, and generatively coherent over time. This approach does not ask whether a model is correct in some external sense, but whether its outputs maintain formal properties, such as constraint propagation or emergent patterning, that make them intelligible and interpretable to humans. In contexts where narrative flow, abstraction, or long-range coherence are key, this kind of structural fidelity may offer insights that token-level accuracy cannot. It opens a path for evaluating systems not just by what they say, but by how they sustain rhythm, pattern, and conceptual balance—qualities that matter both aesthetically and cognitively, even in the absence of explicit meaning.
IV Conclusions
This essay has proposed a structural analogy between large language models and quasicrystals, two systems in which coherence emerges from local constraint without global repetition. Rather than viewing LLMs as imperfect communicators or stochastic approximators, we have suggested understanding them as generators of constrained, non-periodic form. This framing contributes to an ongoing shift in how language and generative systems are conceptualized. In the philosophy of language, it aligns with inferential and structuralist accounts, where meaning arises from relational patterns rather than reference alone. [17] In the philosophy of science, it echoes structural realism, which holds that the epistemic value of scientific theories often lies in the patterns they preserve, not the entities they posit. [18] In AI ethics, it suggests that evaluation itself is not neutral: shifting our metrics toward structural coherence may shape how we justify trust, transparency, and interpretability in generative systems. [19, 21, 20] Just as quasicrystals expanded the scientific definition of order, LLMs may prompt a broader conception of coherence in language and generation.
This perspective invites us to reconsider what it means for language to make sense, and how best to engage with systems that produce such sense without understanding. LLMs do not think in the human sense, but they generate forms marked by pattern, balance, and constraint, qualities we intuitively associate with meaning, even though these systems lack intentionality. The question, then, is not whether such models could give rise to consciousness, but how we might think about them: how to interpret and design around the distinct kind of order they instantiate, and how to recognize the opportunities afforded by their internal structural logic. Like all analogies, this one has boundaries. It does not imply that LLMs share the physical ontology of quasicrystals, nor that linguistic structure is reducible to spatial pattern. Rather, it aims to foreground a neglected mode of evaluation, where structure is coherence without repetition, and to make space for forms of intelligibility that arise without intent. The quasicrystal model serves here not as a literal template, but as a conceptual tool to see coherence where traditional semantics may fall short, and to frame new questions about design, interpretation, and trust in generative systems.